### v2

# Surface-based parcellation and vertex-wise analysis of ultra high-resolution ex vivo 7 tesla MRI in Alzheimer’s disease and related dementias

Abstract
Magnetic resonance imaging (MRI) is the standard modality to understand human brain structure and function in vivo (antemortem). Decades of research in human neuroimaging has led to the widespread development of methods and tools to provide automated volume-based segmentations and surface-based parcellations which help localize brain functions to specialized anatomical regions. Recently ex vivo (postmortem) imaging of the brain has opened-up avenues to study brain structure at sub-millimeter ultra high-resolution revealing details not possible to observe with in vivo MRI. Unfortunately, there has been limited methodological development in ex vivo MRI primarily due to lack of datasets and limited centers with such imaging resources. Therefore, in this work, we present one-of-its-kind dataset of 82 ex vivo T2w whole-brain hemispheres MRI at 0.3 mm3 resolution spanning Alzheimer’s disease and related dementias. We adapted and developed a fast and easy-to-use automated surface-based pipeline to parcellate, for the first time, ultra high-resolution ex vivo brain tissue at the native subject-space resolution using the Desikan-Killiany-Tourville (DKT) brain atlas. This allows us to perform vertex-wise analysis in the template space and thereby link morphometry measures with pathology measurements derived from histology. We open-source our dataset, docker container, Jupyter notebooks for ready-to-use out-of-the-box set of tools and command line options to advance ex vivo MRI clinical brain imaging research at the project webpage.
Keywords:
exvivo MRI 7 tesla surface parcellation segmentation1 Introduction
Alzheimer’s disease and related dementias (ADRD) are heterogeneous, with multiple neuropathological processes jointly contributing to neurodegeneration [34]. Currently, co-pathologies are not reliably detected with in vivo imaging, which makes it difficult for clinicians to accurately diagnose the cause of cognitive impairment in individual patients and to identify those most likely to benefit from treatment with emergent drugs that target AD pathology. The efficacy of in vivo biomarkers can be enhanced by combined analysis of regional measures of neurodegeneration, such as cortical thickness, derived from brain MRI (ex vivo or in vivo) and pathological measures obtained at autopsy. Such analysis can help identify distinct spatial patterns of disease progression linked to AD-specific pathologies (i.e. -amyloid plaques and phosphorylated tau protein tangles) and common co-pathologies. While such studies are performed using either in vivo or ex vivo MRI, the latter offers advantages in temporal proximity to pathological markers and in spatial resolution. The ability to scan ex vivo brain tissue for many hours without motion artifacts makes it possible to achieve ultra-high spatial resolution at which intricate neuroanatomical details of the brain can be revealed [[6], [20], [2], [1], [37] , [32], [36], and [10]]. However, lack of automated tools for processing ultra high-resolution ex vivo MRI is a significant barrier to study structure-pathology association in ADRD.
Decades of neuroimaging research has yielded advanced computational frameworks for automated analysis of in vivo brain MRI, and tools such as FreeSurfer [13] and Statistical Parametric Mapping (SPM) [5], FSL [29] have been applied in a plethora of ADRD neuroimaging studies. However, these tools cannot be directly used on ultra-high resolution ex vivo MRI and there is very limited work on developing ex vivo MRI analysis methods for widespread use, which remains a challenge primarily due to the greater heterogeneity in scanning protocols of ex vivo MRI and increased complexity and imaging artifacts than in vivo MRI. Recent developments for ex vivo MRI include deep learning methods for high resolution cytoarchitectonic mapping in 2D histology [3], atlas-based segmentation of MTL and the thalamus [20], [21], and whole-brain hemispheres analysis [28], [27]. Recently, deep learning-based methods were developed to segment the granular cortical layers [25], and gray matter (GM), subcortical structures, white matter (WM) and its hyperintensities (WMH) in 7T ex vivo MRI. NextBrain [8] provides semi-automatic brain segmentations using histology brain atlas.
These previous methods have several limitations as they are specialized for specific parts of the brain or specific ex vivo datasets mainly limited to a healthy (unremarkable) brain tissue specimens. The methods were evaluated on very small datasets of only 1 [12], 5 [8], and 17 [25] ex vivo brain tissue for unremarkable specimens or at low-resolution [28]. Crucially, to our knowledge, the feasibility of using existing ex vivo MRI analysis pipelines to perform large-scale structure-pathology association studies in ultra-high resolution ex vivo MRI has not been demonstrated.
Contributions. In this study, we perform structure-pathology association analysis in a large dataset of ultra high-resolution 0.3 mm3 T2w 7T MRI scans of whole brain hemispheres from 82 brain donors with ADRD diagnoses, a first study of such scale conducted at this resolution. We present a new computational pipeline that performs automated segmentation and whole-hemisphere FreeSurfer DKT atlas [11] parcellation of the cortex in native subject space at sub-millimeter 0.3 mm 3 resolution, a first large-scale surface-based scheme for ex vivo whole-hemispheres analysis in diseased population. We achieve this by adapting the surface-based modeling steps in FreeSurfer with an initial subject-space topology-corrected WM segmentation derived from a deep learning-based segmentation model as developed in [24]. We evaluate the framework by correlating cortical thickness with neuropathological markers implicated in AD (measures of p-tau, neuronal loss; global amyloid-, Braak staging, and CERAD ratings) and perform vertex-wise generalized linear modeling.
2 Materials and Methods
Dataset. We acquire and analyze a dataset of 82 ex vivo whole hemisphere 0.3 mm3 7T MRI (See Fig. 2) at our XYZ center with left: 42 and right: 40 hemispheres, postmortem interval (PMI) of 18.48 13.60 hours and fixation time of 256.70 280.48 days. The cohort included patients with Alzheimer’s Disease or related dementias (ADRD) [26] spectrum, such as Lewy body disease, limbic-predominant age-related TDP-43 encephalopathy, corticobasal degeneration, primary age-related tauopathy comprising of 41 female (age: 76.97 9.70 years) and 41 male (age: 76.48 11.67 years). Human brain specimens were obtained in accordance with local laws and regulations, and includes informed consent from next of kin at time of death. After autopsy, one hemisphere (Supplemental Fig. 1) was fixed in formalin for atleast 4 weeks and then imaged at T2-w 7T using a 3D-encoded T2 SPACE sequence with 0.3 mm3 isotropic resolution, 3s repetition time (TR), echo time (TE) 383 ms, turbo factor 188, echo train duration 951 ms, bandwidth 348 Hz/px with 2-3 hours per scan.
The non-imaged hemisphere (contralateral tissue) underwent histological processing for neuropathological examination (Supp. Fig. 2). Tissue blocks were embedded in paraffin, sectioned at 6 m thickness, and underwent immunohistochemistry. In each region, semi-quantitative severity ratings of p-tau pathology, amyloid-, and neuronal loss, were assigned by expert neuropathologists on a scale of 0–3 i.e. ‘0: None/Rare’, ‘1: Mild’, ‘2: Moderate’ or ‘3: Severe’. Standardized global AD progression ratings were derived, including Thal stage (global amyloid-), Braak (global p-tau), and CERAD stage (neuritic plaques) [4].
Methodological pipeline. We present a fast and reliable anatomical parcellation and cortical thickness quantification framework that combines deep learning-based methods developed in [24] with classical methods for topology correction [33] and surface-based cortical modeling, inflation, registration, and parcellation in FreeSurfer [[13], [14]] for whole-hemisphere ex vivo brain MRI. Our framework processes one hemisphere in around 60 minutes. Rather than reinventing the wheel, our approach focuses on combining the strengths of these existing tools and making modifications that enable them to scale to ultra-high resolution ex vivo MRI. See Fig. 1 for the schematic pipeline.
Volume-based segmentation: First, we train a deep learning segmentation model nnU-Net [23] as explained in [24] to obtain whole-hemisphere volume-based segmentations of 8 regions namely: cortical gray matter (GM), white matter (WM), white matter hyperintensities (WMH), ventricles and four subcortical structures: caudate, putamen, globus pallidus and thalamus (Fig. 1 column A). The model was trained on Nvidia Quadro RTX 5000 GPU using manual-labeled images and extensively validated, as detailed in the work by Khandelwal et al. [24]. The inference time for each hemisphere is around 15 minutes on a CPU. We merge the WM, WMH, ventricles and subcortical labels in a single label and employ the post-hoc topology correction method [[16] and [7]] based on fast-marching algorithm implemented in ‘Nighres/CRUISE’ [18] to solve the buried sulcus problem, i.e., clearly separating the adjoining gyri in the opposite banks of a buried sulcus. This correction step takes around 5 minutes on a CPU.
Surface-based parcellation: Next, we modify the surface-based scheme in FreeSurfer to obtain DKT atlas-based surface parcellations in ex vivo native subject space. Of note is that this step does not require the raw MRI intensity image and is purely based on the voxel-level segmentation of cortical GM from above. First, a ‘filled’ segmentation with WM, WMH, ventricles and the four subcortical structures as the foreground; and the cortical GM along with the CSF as the background, is created from the initial volume-based segmentation. This ‘filled’ segmentation (Fig. 1 column B) is then tessellated and corrected for any topological errors such as holes and handles based on [33], termed as the ‘WM surface’. The ‘WM surface’ is smoothed and inflated to a sphere (Fig. 1 column C) which is then registered to the spherical representation of the Buckner40 spherical atlas [22] provided in FreeSurfer using the curvature map of the WM surface (Fig. 1 column D). The labels of the DKT-atlas is then warped from the spherical-atlas to the spherical-representation of the ex vivo subject. Next, the ‘WM surface’ is deformed and re-positioned to find the GM-CSF pial surface using the method described in [13] [14] which employs three energy functionals: spring-like term, decomposed into tangential and normal components, to smooth and for regularization of the surface. The third term in the energy functional is dependent on the gradient of the intensity image. This term is the classic stopping criteria (inverse of the image gradient) used in deformable models based on level sets and fast marching methods. FreeSurfer uses raw in vivo T1w (or where applicable T2w) MRI intensity values for this energy term. For the current study, the T2w ex vivo MRI has a varying intensity profile within the GM which causes the deformable surface to prematurely stop in the GM cortical ribbon and does not reach the pial surface. Therefore, we use the segmentation derived from the above volume-based step in place of the T2w ex vivo MRI intensity image. This has a clear advantage as the intensity is constant in the GM cortical ribbon with a zero gradient everywhere. The evolving surface then stops at the pial surface as expected (Fig. 1 column E). Finally, the DKT-atlas labels are projected onto the pial surface and the parcellations (Fig. 1 column F) are mapped to the voxel-space to get the region-wise segmentation in the volume. This surface-based scheme takes 40 minutes on a CPU with 64 threads.
3 Experiments and Results
Segmentation and parcellation. We deployed the deep-learning model to produce the initial segmentation of GM, WM, WMH, ventricles and the subcortical structures. The deep learning-based pipeline was extensively validated in a previous work [24] as explained in Sec. 2. The surface-based scheme was then used to jointly obtain the cortical parcellations and volumetric segmentations as shown in Fig. 2 in the native subject space. Due to the immense manual labelling efforts involved in ex vivo MRI, in the order of 3-4 weeks per subject, the reliability of these parcellations are evaluated using the following two correlation studies.
Region-based cortical thickness vs neuropathology correlations. Fig. 3 shows the Spearman’s -value between the mean cortical thickness (mm) in each brain region (computed in subject space at native resolution) and five pathology measures: global ratings of amyloid-, Braak staging, CERAD, and regional ratings of neuronal loss and p-tau pathology in the MTL, the region first implicated in AD. The analysis were covaried for age, sex and postmortem interval (PMI) and corrected for multiple comparisons using Bonferroni method. Significant negative correlation were found in entorhinal, parahippocampal, medial-orbitofrontal, temporal pole, inferior temporal and parietal lobes which are consistent with literature on progressive loss of cortical gray matter in AD [35].
Surface-based vertex-wise cortical thickness vs neuropathology correlations. The thickness maps for each individual subject were warped to the Buckner40 template-space for vertex-wise correlation analysis between cortical thickness (mm) and the five neuropathological ratings. The parcellations and thickness maps of the three subjects (Fig. 1) in the Buckner40 fsaverage space are shown in Supplemental Fig. 3. Vertex-wise correlation between thickness and the neuropathology ratings was performed by fitting a GLM at each vertex across the entire cohort of 82 subjects with age, sex and PMI as the covariates and corrected for multiple comparisons using family-wise error rate (FWER) correction. Fig. 4 shows the t-statistics map on the pial and the inflated surfaces with the clusters outlined in white indicating regions where the significant strongest associations were observed (p0.05) surviving FDR correction. We observe that the strongest correlations were observed in MTL, the region associated with AD.
Comparison with other methods and parcellations with other atlases. We compared our method with two other segmentation methods SynthSeg [38] (designed for contrast-agnostic MRI) and NextBrain [8] (designed specifically for ex vivo MRI). See Supplemental Fig. 4 for the qualitative results. Both SynthSeg and NextBrain fail to provide meaningful segmentations of ex vivo MRI. SynthSeg (operating at lower 1 mm3) roughly delineates major brain regions but fails at segmenting the finer details in the cortical ribbon and bleeds into CSF. Whereas, NextBrain incorrectly segments the entire cortical ribbon as it does not demarcate the different sulci and gyri. In contrast, our method clearly demarcates the brain regions in native 0.3 mm3 subject-space high resolution using the DKT-atlas. After visual quality control of segmentations produced by both SynthSeg and NextBrain across all the 82 ex vivo subjects, we concluded that a quantitative analysis similar to our proposed method and experiments is not feasible as depicted in Supp. Fig. 4. Separately, in addition to the DKT atlas, we also provide parcellations using the Schaefer [39], Glasser [40] and the Von Economo-Koskinos [41] atlases in native subject-space resolutions for ex vivo brain hemisphere as shown in Supplemental Fig. 5.
4 Discussion
Using both recent advances in deep learning based segmentation models and classic deformable and spherical registration methods, we developed a pipeline that enables surface-based modeling and group-wise registration of the cortical gray matter in ultra high-resolution ex vivo MRI. In addition to its higher resolution, ex vivo MRI poses multiple challenges for computational analysis, including deformation during brain removal and fixation (particularly collapse of the ventricles), and distinct artifacts, such as intensity inhomogeneity both across the whole image domain and across cortical layers. Hence our ability to successfully combine nnU-Net, Nighres/CRUISE and FreeSurfer pipelines to achieve the same type of parcellation as is ubiquitous in in vivo MRI studies is a non-trivial accomplishment. Crucially, the Nighres/CRUISE and FreeSurfer components used in our pipeline operate only on segmentation maps, rather than MRI intensity. Thus, to generalize our approach to other scanners and ultra-high resolution MRI protocols, one only needs to focus on the segmentation aspect.
A key contribution of this work is that it performs, for the first time, a surface-based structure-pathology correlation analysis using ultra-high resolution ex vivo MRI in a large cohort of brain donors with ADRD diagnoses. By doing so, it demonstrates the analysis approach that is ubiquitous in in vivo brain MRI research, and has been previously adopted to resolution ex vivo MRI by some groups [28] and [27], can also be applied to ex vivo studies at ultra-high resolution. To demonstrate analysis modes common in in vivo research, we performed both region of interest (ROI)-level correlation analysis and vertex-wise template-space analysis to study patterns of neurodegeneration within the AD continuum. Prior works [[31], [15]] shows how amyloid- and p-tau etc have differential influences on cortical atrophy. Tau pathology is concurrent with neuronal loss in AD leading to cortical atrophy [17]. Negative correlations between MTL p-tau and neuronal loss with cortical thickness were found to be significant in the entorhinal cortex and the MTL, regions first implicated in AD [[9], [30]] both in in vivo and ex vivo studies. Crucially, the ability to map data from subject space to the DKT-template space based on surface correspondence is not limited to cortical thickness measures, and in future, our pipeline can be used to analyze features that take advantage of ultra-high resolution ex vivo MRI, such as mapping iron and myelin in cortex [36].
Conclusion. We developed a fast and easy-to-use pipeline to parcellate the whole-brain hemispheres into the FreeSurfer DKT-atlas defined regions at the native subject space sub-millimeter 0.3 mm3 isotropic resolution, and applied it to a large-scale dataset of ultra high-resolution ex vivo 7T structural MRI. We demonstrated the feasibility and utility of this approach by performing a structure-pathology association study that is the first of its kind at this level of ex vivo MRI resolution and one of the largest ever conducted at any ex vivo MRI resolution. In future, we will perform an exhaustive quantitative study by delineating the brain regions with different anatomical atlases based on a surface-based population template derived from the ex vivo hemispheres. This will enable us to understand the region-wise local intensity profiles to study the structural relationships with the underlying histology-derived markers. We believe that the present study will enable scientific advancements in ex vivo imaging and thereby inform better in vivo computational biomarkers.
Appendix 0.A Supplementary Figures